# Improving Autonomous AI Agents with Reflective Tree Search and Self-Learning

### v1

@@ Abstract
Autonomous agents have demonstrated significant potential in automating complex multistep decision-making tasks. However, even state-of-the-art vision-language models (VLMs), such as GPT-4o, still fall short of human-level performance, particularly in intricate web environments and long-horizon planning tasks. To address these limitations, we introduce Reflective Monte Carlo Tree Search (R-MCTS), a novel test-time algorithm designed to enhance the ability of AI agents, e.g., powered by GPT-4o, to explore decision space on the fly. R-MCTS extends traditional MCTS by 1) incorporating contrastive reflection, allowing agents to learn from past interactions and dynamically improve their search efficiency; and 2) using multi-agent debate to provide reliable state evaluation. Moreover, we improve the agent’s performance by fine-tuning GPT-4o through self-learning, using R-MCTS generated tree traversals without any human-provided labels. On the challenging VisualWebArena benchmark, our GPT-4o-based R-MCTS agent achieves a 6% to 30% relative improvement across various tasks compared to the previous state-of-the-art. Additionally, we show that the knowledge gained from test-time search can be effectively transferred back to GPT-4o via fine-tuning. The fine-tuned GPT-4o matches 97% of R-MCTS’s performance while reducing compute usage by a factor of four at test time. Furthermore, qualitative results reveal that the fine-tuned GPT-4o model demonstrates the ability to explore the environment, evaluate a state, and backtrack to viable ones when it detects that the current state cannot lead to success. Moreover, our work demonstrates the compute scaling properties in both training - data collection with R-MCTS - and testing time. These results suggest a promising research direction to enhance VLMs’ reasoning and planning capabilities for agentic applications via test-time search and self-learning. [^1] Code and data can be found at https://aka.ms/RMCTS-Self-Learning.
@@ 1 Introduction
In this work, we explore how to effectively scale test-time compute to improve VLM agents, and efficiently transfer knowledge acquired from search back to the model to enhance its reasoning and planning abilities. First, we introduce our Reflective Monte Carlo Tree Search (R-MCTS) agent. Our method extends the classic MCTS with two innovations: contrastive reflection to improve search quality in real time using high-quality reflections obtained from past experiences, inspired by conceptual models of human learning (Marton, 2014); and a multi-agent debate technique for reliable state evaluation, inspired by collaborative and iterative human evaluation approaches (van der Lee et al., 2019). Second, to transfer new knowledge acquired from our R-MCTS agents with increased test-time compute, we propose to update the base VLM with two supervised fine-tuning (SFT) methods, 1) Best-in-Tree SFT, which involves utilizing the optimal, final actions returned from the search process; and 2) Tree-Traversal SFT, which leverages search tree traversals to enable the model to learn how to explore the environment, evaluate states, and backtrack to viable states.
Experiments on the challenging VisualWebArena (Koh et al., 2024a) show that our GPT-4o-based R-MCTS agent achieves a new state-of-the-art performance, with a 6% to 30% relative improvement across various environments compared to the previous state-of-the-art. Additionally, when trained with  Best-in-Tree SFT on only 65 trajectories, GPT-4o can recover 97% of the search performance, while reducing the inference cost by 4x. Furthermore, GPT-4o trained with  Tree-Traversal SFT demonstrates the ability to 1) explore, evaluate, and backtrack without any search augmentation. Moreover, our work yields the compute scaling properties in both training - where data collection with R-MCTS occurs - and testing time, as shown in Figure 1. This represents an important step towards improving VLM’s planning and reasoning capabilities for agentic applications by leveraging test-time search and self-learning.
@@ 3.1 Reflective MCTS
3.2 Self-Learning
Search-augmented agents such as R-MCTS improve performance at the expense of increased test-time computation. Using GPT-4o powered R-MCTS as an example, we explore how to effectively transfer the knowledge gained from tree search back to GPT-4o. We propose two methods: Best-in-Tree SFT (BiT SFT) which directly fine-tunes GPT-4o on the actions returned by the search algorithm, and Tree-Traversal SFT (TrT SFT) which finetunes GPT-4o on all tree traversals that happened during search. We illustrate these two methods in Figure 3.
Given a task $g$, a trajectory $\tau=\{o_{0},a_{0},...,a_{T}\}$, and all search trees $\mathrm{Tree}(o_{0}),..,\mathrm{Tree}(o_{T})$ obtained during R-MCTS, Best-in-Tree SFT trains GPT-4o to learn an improved policy by directly fine-tuning on the actions executed in $\tau$ (i.e., best actions selected by search). This method is simple, but relies on the model itself to learn the implicit reasoning process. Tree-Traversal SFT improves GPT-4o’s decision making ability by using both $\tau$ and tree traversals $\mathrm{Tree}(o_{0}),..,\mathrm{Tree}(o_{T})$. Specifically, given a tree $\mathrm{Tree}(o_{i})$, we 1) replay the tree search process555 This replay stops at the first time when the search process reached the best action. to obtain observations $o$, actions explored or backtracked $a$, and the estimated values $v$; 2) combine value estimation and action into a single action $a\leftarrow(v,a)$, and 3) append them to form a single trajectory $\tau^{\prime}_{i}$. We then repeat this for all trees, and obtain a single trajectory $\tau^{\prime}=\{\tau^{\prime}_{0},...,\tau^{\prime}_{T}\}$ used to train GPT-4o.
Figure 3: Given a task and a trajectory from R-MCTS, Best-in-Tree SFT removes intermediate search trees and directly trains GPT-4o to learn from the final executed actions; Tree-Traversal SFT flattens tree traversals into a single trajectory and trains GPT-4o to explore, evaluate, and backtrack.
@@ 4.1 R-MCTS Results
4.2 Self-Learning Results
@@
React BiT SFT 20 65 2.5x 80.0% 12.4% 31.2%
React TrT SFT 20 35 3.6x 80.0% 18.6% 27.8%
@@
Figure 4: After Tree-Traversal SFT, GPT-4o demonstrates the ability to explore, evaluate, and backtrack without augmenting with search algorithms.
@@
With Tree-Traversal SFT, we find that GPT-4o is able to perform better on unseen tasks by performing exploration, state evaluation, and backtracking without relying on any search algorithm. We illustrate an example in Figure 4: after attempting to directly search for “coffee maker with a touch screen”, the TrT SFT-trained GPT-4o finds that the search function in Classifieds may be unreliable (top search result is a sofa) and backtracked to use category filters instead.
@@ 5.1 Success Rate Breakdown
In addition to having tasks in different environments, VWA labels these task for action difficulty. These are labeled by estimating the number of steps an average human666 Annotations guides are approximate and devised by human annotators of VWA (Koh et al., 2024a). would take to complete the task: easy (1-3 steps), medium (4-9 steps), and hard (10 steps or more). We evaluate all methods using the same setup as in our main results in Section 4.1. We present the results in Table 4.
@@ 7 Conclusion
In this work, we first introduced R-MCTS agent, a search-augmented agent powered by 1) MCTS to explore, evaluate, and backtrack in complex and realistic web environments; 2) a multi-agent-debate value function for reliable state evaluation; and 3) contrastive reflection to learn from past successes and failures in long execution trajectories. Then, we proposed methods to transfer knowledge acquired from search back to GPT-4o by 1) training on best actions returned by tree search (Best-in-Tree SFT); and 2) training on the entire MCTS tree traversal (Tree-Traversal SFT). Experiments on the challenging VisualWebArena benchmark show that R-MCTS achieves a new state of the art, and that our finetuned GPT-4o agent begins to display test-time compute scaling properties similar to search algorithms. We believe our work presents a promising direction to combine test-time search and model self-learning to develop the next-generation autonomous agents.
@@ 8 Limitations
Although MCTS-based agents achieve strong performance on benchmarks such as VWA, they consume nearly 10x more tokens compared to React. This presents a clear trade-off between performance and cost/time, and may limit the practicality of deploying such agents in real-world applications. However, these search-augmented agents can be used to curate training data for VLMs without human labels, which can in turn help develop stronger VLMs. Our work presents a promising step in this direction using Best-in-Tree SFT and Tree-Traversal SFT, and we believe such a self-learning loop is crucial to build stronger agents.

### v2

# Teaching AI Agents to Search with Reflective-MCTS and Exploratory Learning

@@ Abstract
Autonomous agents have demonstrated significant potential in automating complex multistep decision-making tasks. However, even state-of-the-art vision-language models (VLMs), such as GPT-4o, still fall short of human-level performance, particularly in intricate web environments and long-horizon planning tasks. To address these limitations, we present Reflective Monte Carlo Tree Search (R-MCTS) and Exploratory Learning to build o1-like models for agentic applications. We first introduce R-MCTS, a novel test-time algorithm designed to enhance the ability of AI agents to explore decision space on the fly. R-MCTS extends traditional MCTS by 1) incorporating contrastive reflection, allowing agents to learn from past interactions and dynamically improve their search efficiency; and 2) using multi-agent debate to provide reliable state evaluation. Next, we introduce Exploratory Learning, a novel learning strategy to teach agents to search at inference time without relying on any external search algorithms. On the challenging VisualWebArena benchmark, our GPT-4o-based R-MCTS agent achieves a 6% to 30% relative improvement across various tasks compared to the previous state-of-the-art. Additionally, we show that the experience gained from test-time search can be effectively transferred back to GPT-4o via fine-tuning. After Exploratory Learning, GPT-4o 1) demonstrates the ability to explore the environment, evaluate a state, and backtrack to viable ones when it detects that the current state cannot lead to success, and 2) matches 87% of R-MCTS’s performance while using significantly less compute. Notably, our work demonstrates the compute scaling properties in both training - data collection with R-MCTS - and testing time. These results suggest a promising research direction to enhance VLMs’ reasoning and planning capabilities for agentic applications via test-time search and self-learning. [^1] Code and data can be found at https://aka.ms/RMCTS-Self-Learning.
@@ 1 Introduction
In this work, we explore how to effectively scale test-time compute to improve VLM agents, and efficiently transfer experience and knowledge acquired from search back to the model to enhance its reasoning and planning abilities. First, we introduce our Reflective Monte Carlo Tree Search (R-MCTS) agent. Our method extends the classic MCTS with two innovations: contrastive reflection to improve search quality in real time using high-quality reflections obtained from past experiences, inspired by conceptual models of human learning (Marton, 2014); and a multi-agent debate technique for reliable state evaluation, inspired by collaborative and iterative human evaluation approaches (van der Lee et al., 2019). Second, we propose Exploratory Learning, a method to teach agents to search at inference time without relying on any external search algorithms. In contrast to traditional Imitation Learning which trains the model with the optimal, final actions returned from the search process, Exploratory Learning teaches the models to explore the environment, evaluate a state, and backtrack to viable ones using search tree traversals.
Experiments on the challenging VisualWebArena (Koh et al., 2024a) show that our GPT-4o-based R-MCTS agent achieves a new state-of-the-art performance, with a 6% to 30% relative improvement across various environments compared to the previous state-of-the-art. Additionally, we find GPT-4o trained with  Exploratory Learning demonstrates the ability to 1) explore, evaluate, and backtrack without any search augmentation; and 2) recover 87% of the search performance while reducing the inference cost by 2.7x. This represents an important step towards improving VLM’s planning and reasoning capabilities for agentic applications by leveraging test-time search and self-learning.
@@ 3.1 Reflective MCTS
3.2 Exploratory Learning
Search-augmented agents such as R-MCTS improve performance at the expense of increased test-time computation. Using GPT-4o powered R-MCTS as an example, we explore how to teach agents to search at inference time without relying on any external search algorithms. We propose Exploratory Learning (Exploratory), a learning strategy to teach GPT-4o to explore, evaluate, and backtrackon using all tree traversals that happened during search. This is contrast to the traditional Imitation Learning (Imitation) approach, which directly fine-tunes GPT-4o on the actions returned by the search algorithm. We illustrate these two methods in Figure 3.
Given a task $g$, a trajectory $\tau=\{o_{0},a_{0},...,a_{T}\}$, and all search trees $\mathrm{Tree}(o_{0}),..,\mathrm{Tree}(o_{T})$ obtained during R-MCTS, Imitation Learning trains GPT-4o to learn an improved policy by directly fine-tuning on the actions executed in $\tau$ (i.e., best actions selected by search). This method is simple, but relies on the model itself to learn the implicit reasoning process. Exploratory Learning improves GPT-4o’s decision making ability by using both $\tau$ and tree traversals $\mathrm{Tree}(o_{0}),..,\mathrm{Tree}(o_{T})$. Specifically, given a tree $\mathrm{Tree}(o_{i})$, we 1) replay the tree search process555 This replay stops at the first time when the search process reached the best action. to obtain observations $o$, actions explored or backtracked $a$, and the estimated values $v$; 2) combine value estimation and action into a single action $a\leftarrow(v,a)$, and 3) append them to form a single trajectory $\tau^{\prime}_{i}$. We then repeat this for all trees, and obtain a single trajectory $\tau^{\prime}=\{\tau^{\prime}_{0},...,\tau^{\prime}_{T}\}$ used to train GPT-4o.
Figure 3: Given a task and a trajectory from R-MCTS, Imitation Learning removes intermediate search trees and directly trains GPT-4o to learn from the final executed actions; Exploratory Learning flattens tree traversals into a single trajectory and trains GPT-4o to explore, evaluate, and backtrack.
@@ 4.1 R-MCTS Results
4.2 Exploratory Learning Results
@@
React Imitation 20 65 2.5x 80.0% 12.4% 31.2%
React Exploratory 20 35 3.6x 80.0% 18.6% 27.8%
@@
Figure 4: After Exploratory Learning, GPT-4o demonstrates the ability to explore, evaluate, and backtrack without augmenting with search algorithms.
@@
With Exploratory Learning, we find that GPT-4o is able to perform better on unseen tasks by performing exploration, state evaluation, and backtracking without relying on any search algorithm. We illustrate an example in Figure 4: after attempting to directly search for “coffee maker with a touch screen”, the Exploratory Learning trained-GPT-4o finds that the search function in Classifieds may be unreliable (top search result is a sofa) and backtracked to use category filters instead.
@@ 5.1 Success Rate Breakdown
In addition to having tasks in different environments, VWA labels these task for action difficulty. These are labeled by estimating the number of steps an average human777 Annotations guides are approximate and devised by human annotators of VWA (Koh et al., 2024a). would take to complete the task: easy (1-3 steps), medium (4-9 steps), and hard (10 steps or more). We evaluate all methods using the same setup as in our main results in Section 4.1. We present the results in Table 4.
@@ 7 Conclusion
In this work, we first introduced R-MCTS agent, a search-augmented agent powered by 1) MCTS to explore, evaluate, and backtrack in complex and realistic web environments; 2) a multi-agent-debate value function for reliable state evaluation; and 3) contrastive reflection to learn from past successes and failures in long execution trajectories. Then, we proposed methods to transfer knowledge acquired from search back to GPT-4o by 1) training on best actions returned by tree search (Imitation Learning); and 2) training on the entire MCTS tree traversal (Exploratory Learning). Experiments on the challenging VisualWebArena benchmark show that R-MCTS achieves a new state of the art, and that our finetuned GPT-4o agent begins to display test-time compute scaling properties similar to search algorithms. We believe our work presents a promising direction to combine test-time search and model self-learning to develop the next-generation autonomous agents.
@@ 8 Limitations
Although MCTS-based agents achieve strong performance on benchmarks such as VWA, they consume nearly 10x more tokens compared to React. This presents a clear trade-off between performance and cost/time, and may limit the practicality of deploying such agents in real-world applications. However, these search-augmented agents can be used to curate training data for VLMs without human labels, which can in turn help develop stronger VLMs. Our work presents a promising step in this direction using Imitation Learning and Exploratory Learning, and we believe such a self-learning loop is crucial to build stronger agents.